OpenStudy (anonymous):

Can someone help us solve this scientific calculation equation? (5.7 x 10^3)^-4

OpenStudy (hotmama2):

5696?

OpenStudy (anonymous):

rule is (A^n)^m = A^nm for example, squaring a cubed number is equal to the sixth power of the number

OpenStudy (anonymous):

However, your problem is (BxA^n)^m = B^m x A^nm

OpenStudy (anonymous):

(5.7*10^3)^-4 =(5.7)^-4*(10^3)^-4 =1/(5.7)^4*(10)^-12

OpenStudy (anonymous):

x^-4 = (1/x^2)^2 another approach

OpenStudy (ranga):

Express the final answer in scientific notation. That is, it should be of the form: A X 10^B where A is: 1 <= A < 10